OpenStudy (anonymous):

how many tiles 20cm on one side will cover a sala 8m by 6m?

OpenStudy (anonymous):

8x 6 = 48 m squared. 48m = 4800 cm squared Tiles are normally squared so 20 x 20 = 400 so 4800 divided by 400 = ?

OpenStudy (anonymous):

how you get the 4800??

OpenStudy (anonymous):

48m is 4800 cm

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

8m by 6m... that means 8 x6 to find its area...

OpenStudy (campbell_st):

umm a slight error the area is 800 cm by 600 cm so the total area is 480000 cm^2

OpenStudy (anonymous):

Wrong! 48m squared = 48 * (100) ^ 2 cm squared = 480 000 sm squared

OpenStudy (anonymous):

AHh yeah. true

OpenStudy (campbell_st):

so you now you need to find the area of each tile A = 20^2 so the area of a tile is 400 cm^2 number of tiles is 480000/400

OpenStudy (anonymous):

thank you

OpenStudy (campbell_st):

hope it helps